# NEURAL EMBEDDING COMPRESSION FOR EFFICIENT MULTI-TASK EARTH OBSERVATION MODELLING

### v4

@@ 1 INTRODUCTION
Our contributions in this paper are threefold: i) The discussion on EO-FM embedding sharing compared to raw data, ii) the integration of neural compression into the data pipeline for classification and segmentation tasks, and iii) the benchmarking of the neural compression approach on two downstream EO applications.
@@ 3 METHOD
Given an embedding vector $y$, the minimum number of bits required to transmit $y$ is given by Shannon’s source coding theorem [24] as $-\log_{2}p(y)$, where $p$ defines the probability distribution over all embedding vectors. Our goal is to minimize the expectation of this quantity over the distribution of all embedding vectors $-\mathbb{E}\left[\log_{2}p(y)\right]$, which is precisely the entropy of this distribution. We adapt the fully factorized entropy model defined in Ballé et al. [12] in order to model the probability distribution of these embeddings for MAE. Specifically, we leverage the implementation in the CompressAI [25] library.
@@ 4 EXPERIMENTS
In RDC, we compare against transferring the raw data, after entropy coding using the popular Zstandard [27] algorithm, at 16 and 8-bit precision, and fine-tuning the entire model, as shown in Fig. 1 (a). The 16-bit point is a useful reference despite the fact both curated datasets use 8-bit precision since many EO datasets (e.g., Sentinel 2) store data in 16-bit precision.
@@
2. Semantic Segmentation: The Potsdam [31] dataset is used for our semantic segmentation task. This dataset contains large images of urban areas with semantic segmentation labels for six classes. We process it as is standard [32], obtaining tiles of size $512\times 512$. Unlike Wang et al. [3], we keep the ‘clutter’ class and train with labels where the borders of semantic classes are ignored to avoid the ambiguity in labeling these. Additionally, we restrict ourselves to the RGB channels to remain compatible with the MillionAid pre-training. Most state-of-the-art models for semantic segmentation [3, 32, 8] combine embeddings from different stages of the encoder. While NEC can be extended to compress multiple levels of embeddings, we limit our investigation to the single embedding case. We use a simple decoder composed of a first convolutional layer to reduce the embedding dimension by half, followed by six residual convolutional blocks. Then, we use two PixelShuffle [33] blocks to upscale the result by a factor of four. Finally, this output is passed on one branch to a main head with two convolutional layers with $128$ channels and on another branch to an auxiliary head with one convolutional layer with $64$ channels. The loss of the auxiliary head is added to the loss of the main head with a multiplicative factor of $0.4$.
@@ 5 Results
(b) Overall Accuracy on the Potsdam dataset
Fig. 3: Accuracy metric vs Average Bits per Sample for Scene Classification and Semantic Segmentation
@@
1. Scene Classification: Fig. 3(a) plots the results for the UCM Dataset. For the high-performance case, we observe that UQE performs well until $5$ bit quantization and is only slightly outperformed by NEC. For higher compression levels, this does not hold. NEC is able to achieve $94\%$ accuracy at an average embedding size of $0.47$KB, a reduction of only $5\%$ in accuracy for a size reduction of $\sim$700700700700x, greatly outperforming 3 and 2-bit UQE. In particular, transmitting the first layer of the decoder is crucial to retain performance at very high compression.
2. Semantic Segmentation: Results are shown in Fig. 3(b). NEC scales better with compression, achieving more than a $100$x reduction in size with only $5\%$ reduced overall accuracy, compared to a $12\%$ drop in accuracy for comparable compression using UQE. Fig. 4 shows a visual comparison of the inference outputs. For a smaller bit budget on this difficult image, with many separate occurrences of classes to identify, NEC produces more accurate segmentation masks.
In both tasks, we demonstrate the practical viability of our method, with only a small performance gap between RDC and NEC , despite not fine-tuning the encoder.
@@ 6 CONCLUSION
In this work, we introduce NEC, a framework for downstream training and inference for EO tasks based on directly transmitting compressed embeddings. We show how any FM can be adapted to produce compressed embeddings of different fidelities using a single architecture and with minimal compute. Our evaluation demonstrates a data transfer reduction of one to three orders of magnitude with only limited degradation in performance.
We argue the utility of our work from a sustainability point of view, enabling training and inference with limited bandwidth or computational resources and reducing data transfer, storage requirements, and their associated energy costs. Furthermore, we see potential in the long-term storage of compressed embeddings produced from large repositories of EO data (e.g., Landsat, Sentinel) for future analysis.

### v5

@@ 1 INTRODUCTION
Our contributions in this paper are threefold: i) The discussion on EO-FM embedding sharing, ii) the integration of neural compression into the data pipeline for classification and segmentation tasks, and iii) the benchmarking NEC approach on two downstream EO applications.
@@ 3 METHOD
Given an embedding vector $y$, the minimum number of bits required to transmit $y$ is given by Shannon’s source coding theorem [24] as $-\log_{2}p(y)$, where $p$ defines the probability distribution over all embedding vectors. Our goal is to minimize the expectation of this quantity over the distribution of all embedding vectors $-\mathbb{E}\left[\log_{2}p(y)\right]$, which is precisely the entropy of this distribution. We adapt the fully factorized entropy model defined in Ballé et al. [12] in order to model the probability distribution of these embeddings for MAE, leveraging the implementation in the CompressAI [25] library.
@@ 4 EXPERIMENTS
In RDC, we compare against transferring the raw data, after entropy coding using the popular Zstandard [27] algorithm, at 16 and 8-bit precision, and fine-tuning the entire model, as shown in Fig. 1 (a). The 16-bit point is a useful reference despite the fact both curated datasets use 8-bit precision since many EO datasets (e.g., Sentinel 2) store data in greater than 8-bit precision.
@@
Finally, we include JPEG 2000, a widely used and powerful image compression codec. This approach is not entirely comparable to the others, as it aims to transmit the original image, rather than model-ready features, and thus still requires a large encoder to be employed for feature extraction at the consumer. Nonetheless, it provides a good target for performance. For our experiments, we use the original frozen model from Wang et al. [3] as a feature extractor for these images.
@@
2. Semantic Segmentation: The Potsdam [31] dataset is used for our semantic segmentation task. This dataset contains large images of urban areas with semantic segmentation labels for six classes. We process it as is standard [32], obtaining tiles of size $512\times 512$. Unlike Wang et al. [3], we keep the ‘clutter’ class and train with labels where the borders of semantic classes are ignored to avoid the ambiguity in labeling these. Additionally, we restrict ourselves to the RGB channels to remain compatible with the MillionAid pre-training. Most state-of-the-art models for semantic segmentation [3, 32, 8] combine embeddings from different stages of the encoder. While NEC can be extended to compress multi-scale embeddings, we limit our investigation to the single embedding case. We use a simple decoder composed of a first convolutional layer to reduce the embedding dimension by half, followed by six residual convolutional blocks. Then, we use two PixelShuffle [33] blocks to upscale the result by a factor of four. Finally, this output is passed on one branch to a main head with two convolutional layers with $128$ channels and on another branch to an auxiliary head with one convolutional layer with $64$ channels.
@@ 5 Results
(b) Overall accuracy for segmentation on the Potsdam dataset
Fig. 3: Accuracy metric vs bits per pixel. The green line shows uncompressed data and no frozen layers.
@@
1. Scene Classification: Fig. 3(a) plots the results for the UCM Dataset. For the high-performance case, we observe that the simple UQE strategy performs well until $5$ bit quantization (the rightmost point in the plot), however, for higher compression levels, the accuracy rapidly drops. NEC is able to achieve $94\%$ accuracy at an average embedding size of $0.47$KB, a reduction of only $5\%$ in accuracy for a size reduction of $\sim$700700700700x from the raw data, greatly outperforming 3 and 2-bit UQE as well as JPEG2000 in this bpp regime. Retaining the first layer of the decoder is crucial to maintain performance at very high compression. As the rate increases, our method delivers performance comparable to JPEG2000.
2. Semantic Segmentation: Results are shown in Fig. 3(b). NEC scales better than UQE with compression, achieving a $100$x reduction in size with only $5\%$ reduced accuracy, compared to a $12\%$ drop for UQE. Fig. 4 shows a visual comparison of the inference outputs. For a smaller bit budget on this difficult image, with many separate occurrences of classes to identify, NEC produces more accurate segmentation masks. In contrast to scene classification, our method only slightly outperforms JPEG 2000 through the whole bpp range, which may be explained by this task’s greater reliance on high resolution image features. However, JPEG 2000 requires a large encoder to be run as a feature extractor on the client side, whereas NEC directly provides useful features.
@@ 6 CONCLUSION
We introduce NEC, a framework for downstream training and inference for EO tasks based on transmitting compressed embeddings. We show how any FM can be adapted to produce compressed embeddings with minimal compute. We achieve a data transfer reduction of one to three orders of magnitude compared to the original data, with only limited degradation in performance. Additionally, we demonstrate performance comparable to or better than JPEG 2000. We argue the utility of our work from a sustainability point of view, enabling training and inference with limited bandwidth or computational resources as well as long-term storage of data from repositories, thus reducing data transfer, storage requirements, and their associated energy costs. Our codebase will be open-sourced at https://github.com/IBM/neural-embedding-compression/.